OpenStudy (anonymous):

17.Simplify: 5x-5y/5 A.x - 15y B.x + 15y C.x + 3y D.x - 3y 18.Match each example with the appropriate property. A.(5)(1/5) = 1 B.19 + 0 = 19 C.4.2(1) = 4.2 D.8 + -8 = 0 *Additive Inverse *Additive Identity *Multiplicative Inverse *Multiplicative Identity

OpenStudy (anonymous):

You get a medal if you help. (:

OpenStudy (anonymous):

but i dint find any option correct for 17:(

OpenStudy (anonymous):

u sure posted it right?

OpenStudy (anonymous):

18.A is multiplicative inverse B is additive identity C is multiplicative identity D is additive inverse

OpenStudy (anonymous):

Yes I posted it right.

OpenStudy (anonymous):

th e 5 in denominator is under y only?

OpenStudy (anonymous):

No it's suppose to be under the whole thing.

OpenStudy (anonymous):

then its equal t0 x-y,no option is there like this

OpenStudy (angela210793):

17....(5x-5y)/5=5(x-y)/5=x-y